Suppose your travel north for 65 kilometers then travel east 75 kilometers. How far are you from your starting point?
Answer:
Show work here:

Answer by Nate(3500)

.................75..................
....................................
...................................
.................................
...............................
.............................
...........................
.........................
65...................
.....................
...................
..................
................
..............
.............
............
...........
..........
.........
........
.......
......
.....
....
...
..
.
Starting Point ..... A right triangle ....
65^2 + 75^2 = d^2
4,225 + 5,625 = d^2
9,850 = d^2
About 99.25km